OpenStudy (amtran_bus):

Natasha drove from Bedingfield to another city at an average speed of 100 km/h to attend a job interview. On the way back she decided to slow down to enjoy the scenery, so she drove at just 75 km/h. Her trip involved a total of 7 hours of driving time. What is the distance between Bedingfield and the other city?

OpenStudy (amtran_bus):

Any ideas?

ganeshie8 (ganeshie8):

so, distance is same. say \(x\) is the distance between Bedingfield and another city

ganeshie8 (ganeshie8):

use the formula :- distance = speed * time

ganeshie8 (ganeshie8):

while going to interview :- \(x = 100*t_1\) while returning from interviewe :- \(x = 75*(7-t_1)\)

ganeshie8 (ganeshie8):

two equations, two unknowns. you can solve :)

OpenStudy (amtran_bus):

Thank you! I almost did not recognize you since you changed your profile pic!

ganeshie8 (ganeshie8):

lol but i remember the bus :) so u can solve right ? wat do u get for x ?

OpenStudy (amtran_bus):

Is there any way I could set this up as a proportion? But I know x=525-T1 and x=100T1

ganeshie8 (ganeshie8):

yup ! u can set it up as proportion

ganeshie8 (ganeshie8):

but i prefer solving the two equaitons hmm...

OpenStudy (amtran_bus):

Your awesome. Thanks!

ganeshie8 (ganeshie8):

np :) i feel u dont get the method lol -_-

OpenStudy (amtran_bus):

This assignment is for extra credit. It don't count for any type grade.

ganeshie8 (ganeshie8):

ohk good :) u may slowly make peace with solving these then ;)

OpenStudy (amtran_bus):

:)

ganeshie8 (ganeshie8):

problem wid proportion method is, first u need to solve time, and then substitute it in equaiton to get distance. i dont like it that much

ganeshie8 (ganeshie8):

however if the question asks about time, then proporiton method is easy... good luck wid others :)